OpenStudy (anonymous):

I have the answer can someone please check me Where does f(x)=e^x-kx have a relative minimum?

OpenStudy (anonymous):

"relative" so, just the first derivative and second derivatives like before.

OpenStudy (anonymous):

sort of I got x=lnk

OpenStudy (anonymous):

sort of yes.

OpenStudy (anonymous):

good. now, the second derivative would be?

OpenStudy (anonymous):

\[f''(x)=?\]

OpenStudy (anonymous):

no I mean that is the answer for the problem

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[f'(x)=e^x-k=0\\\implies e^x=k\implies x=\ln(k)\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

only one critical point... this is true IF and ONLY if.. k>0

OpenStudy (anonymous):

Im confused so is my answer right or not

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

but.. for reference, "ln" is not defined for numbers "0 or less" so, "k" cannot be "0 or negative"

OpenStudy (anonymous):

that piece of information must be provided with the question.

OpenStudy (anonymous):

well I wrotw the question down exactly

OpenStudy (anonymous):

so, write what I said on your answer and impress your prof. ;)